OpenStudy (anonymous):

A right triangle has one side of length 11 and a hypotenuse of length 12. How long is the third side of the triangle?

OpenStudy (anonymous):

Use the equation:\[a^{2}+b^{2}=c^{2}\] where a and b are the sides and c is the hypotenuse.

OpenStudy (anonymous):

\[Hypotenuse^2 = Base^2 + Height^2\] \[Third Side = \sqrt{144-121} = \sqrt{23}\]